Solution to series-1/2,-1/6,1/6,1/2,5/6,?please send the answer fast thanks

Grade:10

2 Answers

Arun
6 years ago
If we look closely, we get that every number is obtained by adding 2/6 to the preceding number.
As 
-1/2 + 2/6 = -1/6
-1/6 + 2/6 = 1/6
1/6 + 2/6 = 3/6 = 1/ 2
1/2 + 2/6 = 5/6
5/6 + 2/6= 7/6
Hence the next term will be  7/6
 
Akankshya
6 years ago
It is an arithmetic series .Here the common difference is 2/6 ,we can obtain that is we deduct the previous number from the next number.Now if you want the last number then it is 7/6.But if you want their sum then you can apply formula for n terms of sum...sum=n/2 {2a+(n-1)d}..where n is the number of terms,a is the first term,and d is the common difference.